# Short vs. Long-term Coordination of Drones: When Distributed Optimization Meets Deep Reinforcement Learning

### v2

@@ V-A Experimental Settings
Scenario model. In order to evaluate the DO-RL, we model a real-world transportation scenario, where a swarm of drones perform the sensing tasks of traffic monitoring. The scenario involves a real-world map of Munich city, imported to the simulations of urban mobility (SUMO) [^3] https://www.eclipse.org/sumo/ to accurately generate realisitic flows of vehicles. Fig. 3 illustrates a selected map of $1600\times 1600$ meters in the city with the simulation time of $50$ hours, which is $100$ periods. It has a high density of vehicles, approximately $2,000$ vehicles passing by per hour. The area split into a finite number of cells, each is defined as a rectangular square with the size of $200\times 200$ meters that can be captured by the cameras of drones. Fig. 4 shows the distribution of both charging stations and traffic vehicles in the map. The charging stations are uniformly distributed in the map. We employ cross-validation: $80\%$ simulation time of the datasets for training and $20\%$ for testing. There are two scenarios for the experimental evaluation in this paper: (i) Basic scenario. It has $64$ cells, $4$ charging stations and $8$ time periods (set as $30min$ for each period). It has high density of vehicles, around $2,000$ vehicles passing by per time period. $16$ drones sense the area (camera recording) in parallel. The purpose is to understand the influence of different varying parameters in the proposed method. (ii) Complex scenario. It varies the parameter settings such as the density of drones and vehicles, as well as the number of time periods, cells and charging stations. The goal is to identify suitable scenarios for applying the short-term optimization and long-term learning methods and to demonstrate the effective combination of these methods via the proposed DO-RL solution.
@@
Figure 5: DO-RL has its peak overall performance when $J_{u}=2$, $\beta\in[0.1,0.8]$. Change the parameters of the proposed method in mobility range and agents’ behavior.
Neural network and learning algorithm. We sample $H=64$ groups of transitions as minibatches in a replay buffer, and set the discount factor as $0.95$ and the clip interval hyperparameter to $0.2$ for policy updating. Then we try different multi-layer perceptions (MLP), and use $64$ neurons in the three hidden layers of the MLP in both critic and actor networks. The activation function used for the networks is tanh. We train the models for $5000$ episodes.
Complexity. The comparison of both computational and communicational cost is shown in Table III, where $I$ denotes the total number of iterations in EPOS; $|S|$ is the number of states, and $|A|$ is the number of actions.
TABLE III: Comparison of computational and communicational costs.
@@
Computational Cost (agent) $O(L\cdot I)$ $O(|S|\cdot|A|)$ $O(L\cdot I+|S|\cdot|A|$
Computational Cost (system) $O(L\cdot I\cdot\log U)$ $O(|S|\times|A|\cdot U)$ $O((L\cdot I+|S|\cdot|A|)\cdot\log U)$
Communicational Cost (agent) $O(I)$ $O(1)$ $O(I)$
Communicational Cost (system) $O(I\cdot\log U)$ $O(U^{2})$ $O(I\cdot\log U)$
Figure 6: High density of drones increases both mission efficiency and sensing accuracy of all methods, especially for short-term optimization methods (DO-RL and MAPPO). DO-RL also shows its superior performance across different drones densities. Changing the drones density by increasing the number of drones from $8$ to $64$ and fixing 8 periods, 64 cells, 4 charging stations and high density of vehicles.
@@ V-B Baselines and Metrics
MAPPO. It is one state-of-the-art DRL algorithm using PPO [36], but does not include distributed optimization compared to DO-RL. Moreover, agents in MAPPO learn the flying directions timeslot-by-timeslot, rather than period-by-period as in DO-RL. At each timeslot, a drone takes actions to horizontally move to an adjacent cell in eight directions (similar to directions shown in Fig. 1) or hover. Moreover, at the end of every $S$ timeslots (one period), the drones independently identify the closest charging station to recharge. Thus, each agent needs to update the state of its distances to all charging stations.
The evaluation of all algorithms includes three key metrics: 1) Mission efficiency, 2) Sensing accuracy and 3) Energy cost. These metrics are determined using Eq.(1), (2) and (3) respectively. We use the number of vehicles detected/observed by drones to represent the sensor data values defined in this paper. Furthermore, to obtain a comprehensive assessment that considers all three metrics, we conduct an overall performance evaluation using Eq.(4).
@@ V-C Results and Analysis
There are six dimensions for the complex scenario we study here: (i) the density of drones, (ii) the number of periods, (iii) the number of cells, (iv) the number of charging stations, (v) the density of vehicles, (vi) the number of drones and cells while fixing the density of drones, (vii) the remaining battery level, and (viii) the charging load. Results are shown in Fig. 6-11, where the shadow area around the lines represents the standard deviation error of the results.
@@
Density of drones. It denotes the ratio of the number of drones over the number of cells, representing the coverage of a single drone in the map. In this case, we fix the number cells as $64$, and increase the number of drones from $8$ to $64$, that is, the drones density increases from $0.125$ to $1.0$. Fig. 6 illustrates the performance of DO-RL and the three baseline methods when using different drones density to perform traffic monitoring. As drones density increases, both mission efficiency and sensing accuracy of DO-RL increase linearly. These are approximately $23.0\%$ and $15.8\%$ higher than EPOS respectively, as shown in Fig. 6 and Fig. 6. This is because DO-RL controls the traveling direction of drones based on the predicted sensor data, so that drones cover the area with the maximum sensing values. MAPPO has high efficiency and accuracy with a low density of drones (maximum p-value less than $0.001$ using Mann-Whitney U test), but its performance degrades when the drones density is higher than $0.375$ due to the high computational complexity. In contrast, the plan selection in DO-RL improves sensing quality while reducing energy consumption, mitigating the learning difficulty caused by a large number of drones  [^6] DO-RL, EPOS and MAPPO have similar overall performance when the density of drones is $1.0$, with maximum p-value of $0.04$, because there are ample drone resources available to effectively coordinate the coverage of the entire area. . In Fig. 6, the energy cost of DO-RL is on average $1.7\%$ lower than EPOS and $7.9\%$ lower than MAPPO. Greedy has the minimum energy cost among all methods, only $7.89\%$ lower than DO-RL, but has the lowest performance in efficiency and accuracy.
@@
Number of periods. As shown in Fig. 7, DO-RL achieves superior performance compared to other methods, increasing linearly as the number of periods increase (increased by $51.3\%$ in quadruple time). Although it has statistically similar mission efficiency to MAPPO when the number of periods is $4$, this metric increases dramatically because drones update and obtain more accurate predicted sensor data with higher number of periods. The timeslot-by-timeslot learning in MAPPO perplexes the actions learned by drones, particularly considering that it operates with an action space thirty times larger than that of DO-RL, since drones only take actions once per period ($30min$) in DO-RL but take actions per minute in MAPPO. This also results in higher flying energy consumed by drones, and thus the energy cost of MAPPO is around $8.1\%$ higher than DO-RL.
@@
Remaining battery level. It denotes a percentage of battery level once drones have completed their sensing tasks. The value is calculated as an average among all drones within each time period. As shown in the Fig. 10(a) and 10(b), DO-RL keeps drones at a high remaining battery level on high and low density of vehicles, with approximately $30.25\%$ and only $1.05\%$ lower than Greedy. This level follows the drones’ safety regulations which suggest finishing the missions when battery life is around $25\%-30\%$. However, MAPPO does not meet the regulations under low density of vehicles, and the minimum battery level can reach $12.55\%$.
Charging load. It is the total energy demand of drones on each charging station over all time periods. This aims to study the placement of charging stations. As shown in the Fig. 11(a) and 11(b), the total energy demand on charging stations of DO-RL over all periods is more imbalanced compared to other methods. This is because with DO-RL drones learn to travel to the areas with the high required sensing data values. Since drones set the nearest charging stations as destinations, they rely on a lower number of charging places. For example, as shown in Fig. 11(c) and 11(d), when the density of drones is $0.125$ under high density of vehicles, drones only depart from and return to the charging station $D$ since the vehicle distribution around $D$ is more dense than other charging stations. As the density of drones becomes $1$, drones gradually rely on the charging station $B$ and $D$. Similarly, drones only need to fly over the vehicle-dense areas around the charging station $A$ and $B$ under low density of vehicles. These results provide insights to policy makers for building a higher number of charging stations on vehicle-dense areas to support the charging of drones.
In summary, several scientific insights on experimental results are listed as follows: (i) Under scarce drone resources, it is more efficient and accurate to employ long-term learning methods for optimizing sensing. In contrast, under abundant drone resources, short-term optimization methods alone suffice. (ii) In areas with sparse distribution of vehicles over an extended period, long-term learning-based methods prove to be more efficient and accurate compared to short-term optimization. (iii) The proposed method make drones energy-aware and consumes lower energy than traditional DRL, keeping the remaining battery level of drones above the safety line. (iv) The navigation and recharging of drones are limited to the vehicle-dense areas, which recommends policies with lower costs for investments and infrastructure planning by prioritizing the placing of charging stations over these areas. (v) The proposed method effectively combines the strengths of both DRL and distributed optimization to achieve high overall performance under complex sensing environments.
@@ VI Conclusion and Future Work
This paper introduces a new framework named DO-RL to solve the decentralized navigation, sensing and recharging problem by a swarm of drones. DO-RL combines both short-term and long-term strategic coordination; we use distributed optimization to coordinate energy-aware drones to autonomously plan their navigation and sensing within a period, and leverage the deep reinforcement learning to schedule the flying directions of drones. Extensive experimentation using simulations of realistic urban mobility reveals the effectiveness of various settings: Both density of drones and vehicles have a higher influence on the performance of both short-term optimization methods and long-term learning methods compared to the factors such as the number of periods, cells and charging stations. Limited drone resources and sparse vehicle distribution benefit from long-term learning for sensing optimization, while abundant drone resources and high vehicle distribution require only short-term optimization. DO-RL, however, stands out by effectively combining the strengths of both short-term and long-term methods. This combination enables drones to observe traffic flow with higher efficiency and accuracy, even in dynamic and evolving environments. These findings highlight the unprecedented potential of DO-RL in optimizing traffic monitoring and other distributed sensing applications in Smart Cities.

### v3

@@ V-A Experimental Settings
Scenario model. In order to evaluate the DO-RL, we model a real-world transportation scenario, where a swarm of drones perform the sensing tasks of traffic monitoring. The number of vehicles serves as the required sensing value in the model. Drones rely on accurate vehicle observation to detect early signs of traffic congestion. This enables traffic operators to apply mitigation measures, reducing the carbon footprint in one of the world’s highest-emission sectors.. The scenario involves a real-world map of Munich city, imported to the simulations of urban mobility (SUMO) [^3] https://www.eclipse.org/sumo/ to accurately generate realistic flows of vehicles. Fig. 5 illustrates a selected map of $1600\times 1600$ meters in the city with the simulation time of $50$ hours, which is $100$ periods. It has a high density of vehicles, approximately $2,000$ vehicles passing by per hour. The area split into a finite number of cells, each is defined as a rectangular square with the size of $200\times 200$ meters that can be captured by the cameras of drones. Fig. 6 shows the distribution of both charging stations and traffic vehicles in the map. The charging stations are uniformly distributed in the map. We employ cross-validation: $80\%$ simulation time of the datasets for training and $20\%$ for testing. There are two scenarios for the experimental evaluation in this paper: (i) Basic scenario. It has $64$ cells, $4$ charging stations and $8$ time periods (set as $30min$ for each period). It has high density of vehicles, around $2,000$ vehicles passing by per time period. $16$ drones sense the area (camera recording) in parallel. (ii) Complex scenario. It varies the parameter settings such as the density of drones and vehicles, as well as the number of time periods, cells and charging stations.
@@
Figure 7: DO-RL has its peak overall performance when $J_{u}=2$, $\beta\in[0.1,0.8]$. Change the parameters of DO-RL in mobility range and agents’ behavior.
Neural network and learning algorithm. We sample $H=64$ groups of transitions as minibatches in a replay buffer, and set the discount factor as $\gamma=0.95$ and the clip interval hyperparameter to $0.2$ for policy updating. Then we try different multi-layer perceptions (MLP), and use $W=64$ neurons in the three hidden layers of the MLP in both critic and actor networks. The activation function used for the networks is tanh. We train the models for $\mathcal{E}=5000$ episodes with multiple epochs (i.e., time period).
Complexity. Given the number of nodes per layer $W$, number of episodes $\mathcal{E}$, size of state space $\mathcal{X}$, and number of actions $A$, the computational complexity of DNN is approximately $O(C_{dnn}):=O((\mathcal{X}+A)W+3W^{2})$ [18]. The comparison of both computational and communication cost is shown in Table III, where $I$ denotes the number of iterations in EPOS.
TABLE III: Comparison of computational and communication costs.
@@
Computational Cost $O(LI\log U)$ $O(\mathcal{E}TSC_{dnn})$ $O(\mathcal{E}T(C_{dnn}+LI\log U))$
Communication Cost $O(I\log U)$ $O(U^{2})$ $O(I\log U)$
Figure 8: High density of drones increases both mission efficiency and sensing accuracy of all methods, especially for short-term optimization methods (DO-RL and EPOS). DO-RL also shows its superior performance across different drones densities. Changing the drones density by increasing the number of drones from $8$ to $64$ and fixing 8 periods, 64 cells, 4 charging stations and high density of vehicles.
@@ V-B Baselines and Metrics
MAPPO. It is one state-of-the-art DRL algorithm using PPO [35], but does not include distributed optimization compared to DO-RL. Moreover, agents in MAPPO learn the flying directions timeslot by timeslot, rather than period by period as in DO-RL. At each timeslot, a drone takes actions to horizontally move to an adjacent cell in eight directions or hover (similar to the actions in DO-RL), and returns to the nearest charging station at the end of every $S$ timeslots (one period). For fair comparison, this method employs the same reward function and structured tree communication model to share the aggregated observation.
The evaluation of all algorithms includes three key metrics: 1) Mission efficiency, 2) Sensing accuracy and 3) Energy cost. These metrics are determined using Eq.(3), (4) and (5) respectively. We use the number of vehicles detected/observed by drones to represent the sensing values defined in this paper. Furthermore, to obtain a comprehensive assessment that considers all three metrics, we conduct an overall performance evaluation using Eq.(6).
@@ V-C Results and Analysis
There are six dimensions for the complex scenario we study here: (i) the density of drones, (ii) the number of periods, (iii) the number of cells, (iv) the number of charging stations, (v) the density of vehicles, (vi) the number of drones and cells while fixing the density of drones, (vii) the remaining battery level, and (viii) the charging load. Results are shown in Fig. 8-13, where the shadow area around the lines represents the standard deviation error of the results.
@@
Density of drones. It denotes the ratio of the number of drones over the number of cells, representing the coverage of a single drone in the map. In this case, we fix the number cells as $64$, and increase the number of drones from $8$ to $64$, that is, the drones density increases from $0.125$ to $1.0$. Fig. 8 illustrates the performance of DO-RL and the three baseline methods when using different drones density to perform traffic monitoring. As drones density increases, both mission efficiency and sensing accuracy of DO-RL increase linearly. These are approximately $23.0\%$ and $15.8\%$ higher than EPOS respectively, as shown in Fig. 8 and Fig. 8  [^6] DO-RL, EPOS and MAPPO have similar overall performance when the density of drones is $1.0$, with maximum p-value of $0.04$, because there are ample drone resources available to effectively coordinate the coverage of the entire area. . This is because DO-RL controls the traveling direction of drones based on the predicted sensor data, so that drones cover the area with the maximum sensing value. MAPPO has high efficiency and accuracy with a low density of drones (maximum p-value less than $0.001$ using Mann-Whitney U test), but its performance degrades when the drones density is higher than $0.375$ due to the high computational complexity. In contrast, the plan selection in DO-RL reduces the action space and mitigates the learning difficulty, resulting in high performance with a high number of drones. In Fig. 8, the energy cost of DO-RL is on average $1.7\%$ lower than EPOS and $7.9\%$ lower than MAPPO. Greedy has the minimum energy cost among all methods, only $7.89\%$ lower than DO-RL, but has the lowest performance in efficiency and accuracy.
@@
Number of periods. As shown in Fig. 9, DO-RL achieves superior performance compared to other methods, increasing linearly as the number of periods increase (increased by $51.3\%$ in quadruple time). Although it has statistically similar mission efficiency to MAPPO when the number of periods is $4$, this metric increases dramatically because drones update and obtain more accurate predicted sensor data with higher number of periods. The timeslot-by-timeslot learning in MAPPO perplexes the actions learned by drones compared to DO-RL, since drones only take actions once per period ($30min$) in DO-RL but take actions per minute in MAPPO. This also results in higher flying energy consumed by drones, and thus the energy cost of MAPPO is around $8.1\%$ higher than DO-RL.
@@
Remaining battery level. It denotes a percentage of battery level once drones have completed their sensing tasks. The value is calculated as an average among all drones within each time period. As shown in the Fig. 12(a) and 12(b), DO-RL keeps drones at a high remaining battery level on high and low density of vehicles, with approximately $30.25\%$ and only $1.05\%$ lower than Greedy. This level follows the drones’ safety regulations which suggest finishing the missions when battery life is around $25\%-30\%$. However, MAPPO does not meet the regulations under low density of vehicles, and the minimum battery level can reach $12.55\%$.
Charging load. It is the total energy demand of drones on each charging station over all time periods. This aims to study the placement of charging stations. As shown in the Fig. 13(a) and 13(b), the total energy demand on charging stations of DO-RL over all periods is more imbalanced compared to other methods. This is because with DO-RL drones learn to travel to the areas with the high required sensing data values. Since drones set the nearest charging stations as destinations, they rely on a lower number of charging places. For example, as shown in Fig. 13(c) and 13(d), when the density of drones is $0.125$ under high density of vehicles, drones only depart from and return to the charging station $D$ since the vehicle distribution around $D$ is more dense than other charging stations. As the density of drones becomes $1$, drones gradually rely on the charging station $B$ and $D$. Similarly, drones only need to fly over the vehicle-dense areas around the charging station $A$ and $B$ under low density of vehicles. These results provide insights to policy makers for providing higher amount of energy on vehicle-dense areas to support the charging of drones.
In summary, several scientific insights on experimental results are listed as follows: (i) Under scarce drone resources, it is more efficient and accurate to employ long-term learning methods for optimizing sensing. In contrast, under abundant drone resources, short-term optimization methods alone suffice. (ii) Long-term learning methods optimize drone resource usage for traffic management by directing drones towards regions with high vehicle density and advising on infrastructure planning for charging. (iii) With the help of short-term optimization, long-term learning can effectively prevent energy wastage, ensuring that drones maintain their battery levels above safety thresholds.
@@ VI Conclusion and Future Work
This paper introduces a new approach named DO-RL to solve the navigation, sensing and recharging problem by a swarm of drones in a distributed and cost-effective manner. DO-RL leverages the DRL to determine the flying directions of drones period by period, and coordinates energy-aware drones to autonomously plan their navigation and sensing within each period using distributed optimization. Extensive experimentation using simulations of realistic urban mobility reveals the effectiveness of DO-RL: Drones can monitor traffic flow and provide valuable data for optimizing transportation systems and urban planning, ultimately contributing to more efficient and sustainable cybernetic cities. Furthermore, the proposed method effectively optimizes resource utilization, including drone resources, battery safety, and charging. This efficient allocation enables cybernetic systems to scale their drone fleets without significantly increasing resource demands.